# Mittag-Leffler Quantum Statistics and Thermodynamic Anomalies

### v1

@@ VI.2 Evidence for Negative Heat Capacity in Finite Systems
Empirical evidence for negative heat capacity was first reported by D’Agostino et al. [44], who studied multifragmentation in Au + Au collisions at 35 A.MeV. Using event-by-event analysis within the microcanonical ensemble, they reconstructed the excitation energy and temperature, revealing a distinct region of negative heat capacity around 5 A.MeV. This observation aligns with theoretical predictions that finite systems undergoing a first-order phase transition may develop a convex entropy branch and exhibit negative microcanonical heat capacity.
Further support was provided by D’Agostino et al. [45] through a detailed analysis of caloric curves and energy fluctuations during the liquid-gas phase transition in finite nuclear systems. Their results demonstrated a clear thermodynamic signature in the microcanonical caloric curve, highlighting the necessity of microcanonical analysis for capturing such noncanonical behavior.
@@
On the theoretical side, Padmanabhan [47] showed that self-gravitating systems with long-range interactions inherently exhibit negative heat capacity within the microcanonical ensemble. Employing toy models and mean-field approximations, the study revealed that these systems transition between kinetic-energy-dominated and potential-energy-dominated phases through an intermediate regime characterized by negative heat capacity. This regime corresponds to virial equilibrium and disappears in the canonical ensemble, where it is replaced by a discontinuous phase transition. These findings emphasize the inequivalence of statistical ensembles and the central role of the microcanonical framework in describing systems with long-range interactions.
Additional theoretical support comes from the microcanonical formalism developed by Gross [48], which investigates phase transitions in small systems beyond the thermodynamic limit. It is argued that the microcanonical heat capacity can be extracted from fluctuations in the kinetic energy of nuclear fragments. In multifragmenting nuclei, large fluctuations in kinetic energies, combined with a convex entropy curvature, lead to negative microcanonical heat capacity values.
@@ VII.1 Explanation of the Tsallis Function
The Tsallis function  [49, 31], or $q$-entropy, is a generalization of the Boltzmann-Gibbs (BG) entropy introduced by Tsallis to describe systems with nonextensive properties. It is defined as:
@@ VII.2 Explanation of the Kaniadakis Function
The Kaniadakis function [50, 32], or $\kappa$-exponential, is a one-parameter deformation of the standard exponential function, introduced by Kaniadakis to generalize BG statistics. It is defined as:
@@ VII.3 Thermodynamic Geometry
In the Tsallis framework  [31], the thermodynamic curvature exhibits behavior largely consistent with the standard statistics. For Tsallis-BE with $q>1$, the curvature is positive, indicating attractive interactions, and diverges at fugacity $z=1$, a signature of BE condensation. Increasing $q$ reduces the magnitude of $R$, suggesting enhanced stability. In contrast, Tsallis-BE gases display negative curvature, consistent with repulsive interactions, while nonextensive Tsallis-BE gases remain interaction-free, with $R=0$. Thus, the Tsallis parameter modifies the intensity of statistical interactions but does not alter their qualitative nature.
By comparison, the Kaniadakis framework  [32] reveals more nuanced geometric behavior. In $\kappa$-deformed MB systems, $R$ deviates from the classical zero value and becomes negative, indicating $\kappa$-induced repulsion. $\kappa$-deformed Fermi systems consistently exhibit negative curvature, as expected from fermionic repulsion. For $\kappa$-deformed Bose systems, however, the curvature is more intricate: below a critical fugacity $z^{*}$, $R$ is negative, resembling fermionic interactions, while above $z^{*}$, it becomes positive, reflecting bosonic attraction. At $z=1$, the divergence of $R$ signals BE condensation.
@@
The central point is that, as demonstrated in Reference  [11], the MLMB statistic —a classical ML-based statistic— exhibits distinct interaction behaviors that are contingent upon the value of the parameter $\alpha$. Specifically, for $\alpha>1$, the interactions are attractive, whereas for $\alpha<1$, they are repulsive. When $\alpha=1$, the results recover the classical MB statistics, corresponding to a non-interacting system. These findings suggest that MLMB statistics inherently encompass both attractive and repulsive interactions, in contrast to the classical Tsallis statistic, which operates under a purely non-interacting framework, and the classical Kaniadakis statistic, which presumes solely repulsive thermodynamic curvature.
@@ VII.4 Heat Capacity
In the context of classical MLMB statistics, as demonstrated in [11], the heat capacity reduces to the standard classical value, $C_{V}=3/2$, when $\alpha=1$. For $\alpha>1$, the system exhibits behavior analogous to the MLBE case with $\alpha>1$, while for $\alpha<1$, the ratio $C_{V}/(Nk_{B})$ diverges at low temperatures and asymptotically approaches zero at higher temperatures, without ever reaching it exactly.
@@ VIII The MLBE statistics Debye solid
Following the approach presented in [51], we adopt the Debye framework. In the present work, we extend this formulation by replacing the conventional BE statistics with the generalized MLBE statistics.
@@
The curves in the  LABEL:CVBE_Debye diagram show that the heat capacity begins at zero and increases monotonically with temperature. At low temperatures, this increase is relatively steep, but the growth rate diminishes at higher temperatures until the heat capacity approaches an asymptotic constant value. The case of $\alpha=1$ corresponds to the standard case. For $\alpha<1$, the heat capacity remains consistently below the standard Debye model, yielding curves positioned beneath the $\alpha=1$ across the entire temperature range. In contrast, for $\alpha>1$, the heat capacity exceeds the standard case, resulting in curves that lie above the case of $\alpha=1$. This classification highlights the role of the parameter $\alpha$ in shifting the magnitude of the heat capacity relative to the standard Debye model.
@@
    Table 1 presents the corresponding values of $C_{V}$ in the high-temperature regime for different values of $\alpha$. These values are in agreement with the trends observed in Figure LABEL:CVBE_Debye, which is generated from equation (27).
@@
    Figure LABEL:LOWER_ML1 depicts the variation of constant-volume heat capacity within the Debye model across different values of the deformation parameter $\alpha$. The heat capacity demonstrates a linear relationship with the normalized temperature $(T/\theta)^{3}$, with an increasing slope of the curves corresponding to higher values of $\alpha$.
@@ VIII.1 Generalized Debye Model with $q$-Deformed
The thermodynamic anomalies identified in the MLBE and MLFD frameworks—namely, the anomalous high-temperature heat capacity in the MLBE case and the emergence of negative heat capacity in the MLFD scenario—are not phenomena unique to ML-based statistics. Analogous features have been documented within other generalized statistical formalisms, particularly those grounded in $q$-deformed quantum statistics derived from the Tamm–Dancoff (TD) oscillator algebra. In recent work, Chung and Algin  [52] examined the thermodynamic properties of a $q$-deformed Debye model, wherein phonons are treated as excitations obeying a TD-type $q$-deformed bosonic algebra.
@@
They obtain the heat capacity for the q-deformed Debye solid as follows [52]
@@ VIII.2 Comparative Analysis
The preceding sections established the heat capacity within the Debye framework generalized using the MLBE statistics. In parallel, $q$-deformed Debye models [52] provide an alternative generalized-statistics formulation in which phonons obey a TD-type $q$-deformed bosonic algebra.
@@
    which for $\alpha=1$ recovers the classical Dulong–Petit value $C_{V}\rightarrow 3Nk_{B}$. In contrast, the $q$-deformed Debye solid exhibits [52]
@@
    where the polylogarithm function $\mathrm{Li}_{4}(q)$ is defined by the series $\mathrm{Li}_{s}(z)=\sum_{k=1}^{\infty}z^{k}/k^{s}$ [52].

### v2

@@ VI.2 Evidence for Negative Heat Capacity in Finite Systems
Empirical evidence for negative heat capacity was first reported by D’Agostino et al. [d1999thermodynamical], who studied multifragmentation in Au + Au collisions at 35 A.MeV. Using event-by-event analysis within the microcanonical ensemble, they reconstructed the excitation energy and temperature, revealing a distinct region of negative heat capacity around 5 A.MeV. This observation aligns with theoretical predictions that finite systems undergoing a first-order phase transition may develop a convex entropy branch and exhibit negative microcanonical heat capacity.
Further support was provided by D’Agostino et al. [d2000negative] through a detailed analysis of caloric curves and energy fluctuations during the liquid-gas phase transition in finite nuclear systems. Their results demonstrated a clear thermodynamic signature in the microcanonical caloric curve, highlighting the necessity of microcanonical analysis for capturing such noncanonical behavior.
@@
On the theoretical side, Padmanabhan [padmanabhan1990statistical] showed that self-gravitating systems with long-range interactions inherently exhibit negative heat capacity within the microcanonical ensemble. Employing toy models and mean-field approximations, the study revealed that these systems transition between kinetic-energy-dominated and potential-energy-dominated phases through an intermediate regime characterized by negative heat capacity. This regime corresponds to virial equilibrium and disappears in the canonical ensemble, where it is replaced by a discontinuous phase transition. These findings emphasize the inequivalence of statistical ensembles and the central role of the microcanonical framework in describing systems with long-range interactions.
Additional theoretical support comes from the microcanonical formalism developed by Gross [gross2001microcanonical], which investigates phase transitions in small systems beyond the thermodynamic limit. It is argued that the microcanonical heat capacity can be extracted from fluctuations in the kinetic energy of nuclear fragments. In multifragmenting nuclei, large fluctuations in kinetic energies, combined with a convex entropy curvature, lead to negative microcanonical heat capacity values.
@@ VII.1 Explanation of the Tsallis Function
The Tsallis function  [Tsallis1988possible, adli2019nonperturbative], or $q$-entropy, is a generalization of the Boltzmann-Gibbs (BG) entropy introduced by Tsallis to describe systems with nonextensive properties. It is defined as:
@@ VII.2 Explanation of the Kaniadakis Function
The Kaniadakis function [kaniadakis2001non, mehri2020thermodynamic], or $\kappa$-exponential, is a one-parameter deformation of the standard exponential function, introduced by Kaniadakis to generalize BG statistics. It is defined as:
@@ VII.3 Thermodynamic Geometry
In the Tsallis framework  [adli2019nonperturbative], the thermodynamic curvature exhibits behavior largely consistent with the standard statistics. For Tsallis-BE with $q>1$, the curvature is positive, indicating attractive interactions, and diverges at fugacity $z=1$, a signature of BE condensation. Increasing $q$ reduces the magnitude of $R$, suggesting enhanced stability. In contrast, Tsallis-BE gases display negative curvature, consistent with repulsive interactions, while nonextensive Tsallis-BE gases remain interaction-free, with $R=0$. Thus, the Tsallis parameter modifies the intensity of statistical interactions but does not alter their qualitative nature.
By comparison, the Kaniadakis framework  [mehri2020thermodynamic] reveals more nuanced geometric behavior. In $\kappa$-deformed MB systems, $R$ deviates from the classical zero value and becomes negative, indicating $\kappa$-induced repulsion. $\kappa$-deformed Fermi systems consistently exhibit negative curvature, as expected from fermionic repulsion. For $\kappa$-deformed Bose systems, however, the curvature is more intricate: below a critical fugacity $z^{*}$, $R$ is negative, resembling fermionic interactions, while above $z^{*}$, it becomes positive, reflecting bosonic attraction. At $z=1$, the divergence of $R$ signals BE condensation.
@@
The central point is that, as demonstrated in Reference  [seifi2025intrinsic], the MLMB statistic —a classical ML-based statistic— exhibits distinct interaction behaviors that are contingent upon the value of the parameter $\alpha$. Specifically, for $\alpha>1$, the interactions are attractive, whereas for $\alpha<1$, they are repulsive. When $\alpha=1$, the results recover the classical MB statistics, corresponding to a non-interacting system. These findings suggest that MLMB statistics inherently encompass both attractive and repulsive interactions, in contrast to the classical Tsallis statistic, which operates under a purely non-interacting framework, and the classical Kaniadakis statistic, which presumes solely repulsive thermodynamic curvature.
@@ VII.4 Heat Capacity
In the context of classical MLMB statistics, as demonstrated in [seifi2025intrinsic], the heat capacity reduces to the standard classical value, $C_{V}=3/2$, when $\alpha=1$. For $\alpha>1$, the system exhibits behavior analogous to the MLBE case with $\alpha>1$, while for $\alpha<1$, the ratio $C_{V}/(Nk_{B})$ diverges at low temperatures and asymptotically approaches zero at higher temperatures, without ever reaching it exactly.
@@ VIII The MLBE statistics Debye solid
Following the approach presented in [kittel2018introduction], we adopt the Debye framework. In the present work, we extend this formulation by replacing the conventional BE statistics with the generalized MLBE statistics.
@@
The curves in the Fig.  9 diagram show that the heat capacity begins at zero and increases monotonically with temperature. At low temperatures, this increase is relatively steep, but the growth rate diminishes at higher temperatures until the heat capacity approaches an asymptotic constant value. The case of $\alpha=1$ corresponds to the standard case. For $\alpha<1$, the heat capacity remains consistently below the standard Debye model, yielding curves positioned beneath the $\alpha=1$ across the entire temperature range. In contrast, for $\alpha>1$, the heat capacity exceeds the standard case, resulting in curves that lie above the case of $\alpha=1$. This classification highlights the role of the parameter $\alpha$ in shifting the magnitude of the heat capacity relative to the standard Debye model.
@@
    Table 1 presents the corresponding values of $C_{V}$ in the high-temperature regime for different values of $\alpha$. These values are in agreement with the trends observed in Figure 9, which is generated from equation (27).
@@
    Figure 10 depicts the variation of constant-volume heat capacity within the Debye model across different values of the deformation parameter $\alpha$. The heat capacity demonstrates a linear relationship with the normalized temperature $(T/\theta)^{3}$, with an increasing slope of the curves corresponding to higher values of $\alpha$.
@@ VIII.1 Generalized Debye Model with $q$-Deformed
The thermodynamic anomalies identified in the MLBE and MLFD frameworks—namely, the anomalous high-temperature heat capacity in the MLBE case and the emergence of negative heat capacity in the MLFD scenario—are not phenomena unique to ML-based statistics. Analogous features have been documented within other generalized statistical formalisms, particularly those grounded in $q$-deformed quantum statistics derived from the Tamm–Dancoff (TD) oscillator algebra. In recent work, Chung and Algin  [chung2024q] examined the thermodynamic properties of a $q$-deformed Debye model, wherein phonons are treated as excitations obeying a TD-type $q$-deformed bosonic algebra.
@@
They obtain the heat capacity for the q-deformed Debye solid as follows [chung2024q]
@@ VIII.2 Comparative Analysis
The preceding sections established the heat capacity within the Debye framework generalized using the MLBE statistics. In parallel, $q$-deformed Debye models [chung2024q] provide an alternative generalized-statistics formulation in which phonons obey a TD-type $q$-deformed bosonic algebra.
@@
    which for $\alpha=1$ recovers the classical Dulong–Petit value $C_{V}\rightarrow 3Nk_{B}$. In contrast, the $q$-deformed Debye solid exhibits [chung2024q]
@@
    where the polylogarithm function $\mathrm{Li}_{4}(q)$ is defined by the series $\mathrm{Li}_{s}(z)=\sum_{k=1}^{\infty}z^{k}/k^{s}$ [chung2024q].